Vatic Capital. Cachita Haynes works as a currency speculator for Vatic Capital of Los Angeles. Her latest speculative position is to profit from her expectation that the U.S. dollar will rise significantly against the Japanese yen. The current spot rate is ​¥119.00​/$. She must choose between the following 90​-day options on the Japanese​ yen: LOADING.... a. Should Cachita buy a put on yen or a call on​ yen? b. What is​ Cachita's breakeven price on the option purchased in part a​? c. Using your answer from part a​, what is​ Cachita's gross profit and net profit​ (including premium) if the spot rate at the end of 90 days is ​¥140.00​/$? Option Strike Price Premium Put on yen yen128​/$. $0.00004​/yen Call on yen yen128​/$ $0.00049​/yen
Answer #1

Solution:

1) Cachita should be buying a put on Yen as she expects the USG to rise or Yen to depreciate so Yen Put is what she would be looking for.

2) Breakeven price would be :

Strike Price + Premium paid = 128 + .00004*119 = 128 + .00476 = 128.00476 Yen

3) Cachita gross profit = Price at the expiry - Strike Price =140 - 128 = 12 yen

Net Profit = Price at expiry - Strike - Premium Paid = 140 - 128 - .00476 = 12 - .00476 = 11.99524 Yen